# Federated Learning Clients Clustering with Adaptation to Data Drifts

### v1

@@
Abstract.
Federated Learning (FL) enables deep learning model training across edge devices and protects user privacy by retaining raw data locally. Data heterogeneity in client distributions slows model convergence and leads to plateauing with reduced precision. Clustered FL solutions address this by grouping clients with statistically similar data and training models for each cluster. However, maintaining consistent client similarity within each group becomes challenging when data drifts occur, significantly impacting model accuracy. In this paper, we introduce Fielding, a clustered FL framework that handles data drifts promptly with low overheads. Fielding detects drifts on all clients and performs selective label distribution-based re-clustering to balance cluster optimality and model performance, remaining robust to malicious clients and varied heterogeneity degrees. Our evaluations show that Fielding improves model final accuracy by 1.9%-5.9% and reaches target accuracies 1.16$\times$-2.61$\times$ faster.
copyright: nonedoi: isbn: price:
1. Introduction
@@
Federated Learning (FL) is a distributed machine learning paradigm that enables multiple clients to train one model collaboratively while retaining raw data locally (Bonawitz et al., 2019; McMahan et al., 2017). As FL circumvents the communication costs of data movement and mitigates privacy leakage (Lyu et al., 2024; Huba et al., 2022), it has drawn attention across various fields and seen adoption in applications such as Google Keyboard (Yang et al., 2018), energy consumption forecasting (Abdulla et al., 2024), and medical image processing (Zhou et al., 2023).
The main challenge in federated learning is data heterogeneity. Since federated learning involves many devices of diverse settings, there exists substantial dataset size and distribution divergences between clients from distant geographical locations and with distinct personal habits (Li et al., 2022a; Kim et al., 2024). Such data heterogeneity leads to slow model convergence and plateauing with degraded precision (Li et al., 2020a; Zhao et al., 2018; Sattler et al., 2020). A wide range of clustering mechanisms have been proposed to group clients with similar data distribution into more homogeneous clusters and train one model for each cluster (Ghosh et al., 2020; Duan et al., 2021; Liu et al., 2023; Jothimurugesan et al., 2023).
Data drift occurs frequently in federated learning systems. It is defined as temporal changes in the label distribution (i.e., the proportion of samples assigned to each label) of clients’ local datasets. Such drifts arise from varying sampling rates, hardware configurations, sensor modalities, and limitations in client storage capacities (Chen et al., 2020; Nandi and Xhafa, 2022; Marfoq et al., 2023). Existing clustering mechanisms, however, are susceptible to data drift. Prior works address this issue but often either inadequately adjust clusters or incur prohibitive costs. IFCA and FlexCFL (Ghosh et al., 2020; Duan et al., 2021) adjust clusters by relocating shifted clients while maintaining a fixed number of clusters, operating under the assumption of low drift magnitude. Auxo (Liu et al., 2023) re-clusters clients participating in the current training round using gradient information; however, since only a subset of clients is selected each round, it delays adjustments for drifted clients until they are chosen for training. FedDrift (Jothimurugesan et al., 2023) re-clusters all clients and adaptively determines the number of clusters by transmitting all cluster models to every client each round to obtain training loss on local data, leading to significant communication and computation costs.
In this work, we propose Fielding, a novel clustering-based FL framework that retains cluster client similarity despite frequent data drifts. Fielding focuses on re-clustering clients as data drifts, dynamically adjusting the number of clusters while stabilizing responses to minor data drifts. Fielding combines per-client migration with selective global clustering to capture significant distribution patterns without compromising accuracy.
We assume the existence of a coordinator, typically deployed on a powerful central server to manage synchronization and advance global model training (Bonawitz et al., 2019; Bagdasaryan et al., 2020). Our solution handles all drifted clients by clustering them according to their local distribution vector, a piece of information readily available all the time. We present a theoretical framework supporting this approach. Fielding is designed to be lightweight; it introduces no additional computation on the client side and requires clients only to transmit their distribution vector (an array with up to a few thousand values) when they register with the coordinator.
We implement Fielding by extending the widely adopted FL engine FedScale (Lai et al., 2022) to support streaming data and our clustering mechanisms. We evaluated Fielding with up to 5078 clients on four image streaming traces with data drifts. Compared to existing approaches, Fielding improves model accuracy by 1.9% to 5.9% and achieves the target accuracy 1.16$\times$ to 2.61$\times$ faster. We also show that Fielding is compatible with various client selection and aggregation algorithms, and robust to malicious clients and varied heterogeneity degrees.
Our main contributions are:
  • We introduce Fielding, a clustered FL framework that handles data drifts promptly with low overheads.
  • We show that Fielding imposes minimal overhead by eliminating additional client-side computations and reducing communication costs through the use of readily available data, and we provide a theoretical framework to support our approach.
  • We extend the FedScale FL engine to implement Fielding and demonstrate its efficacy on four real-world datasets with up to 5,078 clients, achieving up to 5.9% accuracy improvement and faster convergence compared to existing methods.
  • We show that Fielding is compatible with various client selection and aggregation algorithms and is robust against malicious clients and different degrees of data heterogeneity.
2. Background
In this section, we provide the background of clustered federated learning and the impacts of data drifts.
2.1. Clustered Federated Learning
FL deployment often involves hundreds to thousands of available clients participating in the training (Bonawitz et al., 2019). For example, Google sets 100 as the target number of clients for a round of training when improving keyboard search suggestions (Yang et al., 2018) and (Huang et al., 2019) have developed mortality and stay time predictor for the eICU collaborative research database (Pollard et al., 2018) containing data of 208 hospitals and more than 200,000 patients. Although thousands of devices may be available in a given round, FL systems typically select a subset to participate in training. This selection process ensures a reasonable training time and accounts for diminishing returns where including more clients doesn’t accelerate convergence (Xu et al., 2024).
Data heterogeneity is a major challenge in FL. Varying volumes and distribution of local data among participants is known to cause slow convergence and poor model accuracy (Yang et al., 2021; Wen et al., 2023). To mitigate the impact of heterogeneity, an effective approach is clustering, which groups statistically similar clients. Clients in the same cluster train one cluster model collaboratively and also have their inference requests served by this model. Previous works such as Auxo (Liu et al., 2023), FlexCFL (Duan et al., 2021), FedDrift (Jothimurugesan et al., 2023), and FedAC (Zhang et al., 2024) have shown that clustering can achieve high accuracy and fast model convergence.
2.2. Data Drifts Affect Clustering
The key challenges for clustered federated learning are data drifts. Data drifts refer to the changes in the percentage of clients’ local samples associated with a specific label. Drifts happen naturally when clients have access to non-stationary streaming data (e.g., cameras on vehicles and virtual keyboards on phones) (Koh et al., 2021; Bhardwaj et al., 2022). For example, when training an FL model for keyboard suggestions, drifts occur when many clients simultaneously start searching for a recent event (widespread drift), a few clients pick up a niche hobby (concentrated intense drift), or a new term appears when a product is launched (new label added).
Data drifts make clustered FL ineffective as it increases the data heterogeneity in the clusters. We demonstrate this problem with the Functional Map of the World (FMoW) dataset (Christie et al., 2018). This dataset contains time-stamped satellite images labeled according to the function (e.g., recreational facility, crop field, etc.) of the captured land. We keep images taken on or after January 1, 2015, and create one client per unique UTM zone metadata value (302 clients in total). A day’s worth of data becomes available every two training rounds, and clients maintain images they received over the last 100 rounds.
We use mean client distance as the intra-cluster heterogeneity metric (Lai et al., 2021). For each client, we calculate the average pairwise L1 distance between its data distribution vector and those of all other clients in the same cluster. We then use the mean of these per-client distances to quantify the overall heterogeneity after clustering. Note that instead of first averaging within each cluster and then finding the mean of those cluster averages, we use the overall mean across all clients. This helps us avoid biased results arising from imbalanced cluster sizes. Note that in the baseline case without any clustering, we consider all clients as members of a “global cluster”.
@@ 2.2. Data Drifts Affect Clustering
Figure 1. Client heterogeneity of the global set and clusters generated through label-based clustering.
We cluster clients based on their label distribution vectors at the starting round and track the overall heterogeneity by measuring the mean client distance as described earlier. We cluster based on label distribution because data drifts occurs when label distribution changes. Figure 1 shows that at the beginning (e.g., round 1), clustering reduces the per-cluster heterogeneity compared with no clustering (i.e., putting all clients in a global set). However, as data distribution shifts over time, static clustering increases per-cluster heterogeneity and soon gets close to the no clustering case at round 322.
3. Fielding Design decisions
Addressing data drifts is crucial for preserving the similarity of clients within a cluster. To handle data drifts, we must identify which clients should be moved, determine the appropriate clusters for these clients, and specify the information to collect for drift detection and client movements.
3.1. Combine Per-client Adjustment and Re-clustering
Existing clustered FL works rely on per-client adjustment or re-clustering to handle data drifts. However, neither approach works well on all levels of drifts.
Per-client adjustment fails in many-drift scenarios. Per-client adjustment keeps a fixed number of clusters and moves drifted clients individually to another cluster with similar gradients or model weights (Ghosh et al., 2020; Duan et al., 2021). However, in highly dynamic environments where many clients change their distributions at the same time (Jothimurugesan et al., 2023), we have to move all these clients across clusters and the cluster membership changes dramatically. It is hard to decide which cluster the client should move to because the basis used by previous works to match client to the closest cluster (i.e., cluster’s average model weights or gradients) is changing (Ghosh et al., 2020; Duan et al., 2021). We demonstrate this problem in Figure 1 (“Move Individuals”). In rounds 1-40 and rounds 410-615, individual movements help keep the mean client distance below that of the global set even though clients are becoming more heterogeneous (signaled by the increasing heterogeneity value). However, as the global set of all clients experiences major drifts during rounds 600-730 (signaled by the heterogeneity spike), many clients start moving across clusters simultaneously. Per-client adjustment then results in poor clusters and an overall heterogeneity value higher than the global set in rounds 740-1340.
Moreover, maintaining a fixed number of clusters can lead to sub-optimal clustering when new distribution patterns emerge during data drifts. For instance, we want to create another cluster when a new label appears for clients with a substantial number of samples of that label, but keeping the cluster number static will deem such creation infeasible.
(a) Selective global re-clustering vs. default global re-clustering. Positive numbers mean selective re-clustering is better.
(b) Global re-clustering vs. Selected re-clustering. Positive numbers mean re-clustering all drifted clients is better.
Figure 2. Accuracy difference between different re-clustering approaches.
Re-clustering makes model unstable during small drifts In Figure 1, we show an ideal case where we get optimal clusters by globally re-clustering all clients using label distribution vectors when any client drifts (“Global recluster”). Our implementation bounds the maximum number of clusters to 10, as previous works empirically show that having more clusters can negatively affect model convergence due to limited training resources (Liu et al., 2023). As expected, this approach yields the lowest heterogeneity value. However, we will next show that such low heterogeneity does not necessarily lead to superior model accuracy.
Common clustering algorithms such as k-means have randomized initialization, leading to distinct results in multiple runs (Arthur and Vassilvitskii, 2006; Ahmed et al., 2020). This behavior is also observed in our experiment. In Figure 1, the “Global Recluster” curve exhibits more dramatic changes between rounds than other curves. Following the global re-clustering process, we need to determine the model each cluster should continue with. In our implementation, we align with previous works and use the average of the clients’ models within each cluster (Long et al., 2023; Zhang et al., 2024). When most clients remain in the same cluster, the new cluster models would be similar to the previous ones. But when there are significant changes in the clusters, either due to large drifts or bad centroid initialization, the new models will differ substantially and require additional time to stabilize. This leads to sudden accuracy drops as the new clusters warm up. In Figure 2(a), the x-axis denotes training rounds and the y-axis shows the mean accuracy we get with our selective global re-clustering (details in Section 3) subtracted by that of triggering global re-clustering by default. We observe that default global re-clustering leads to fluctuating test accuracy, which oscillates between meeting our approach’s accuracy and being as much as 5% worse.
Summary. Per-client adjustment fails in highly dynamic environments where many clients experience drifts. In contrast, global re-clustering may harm model accuracy because it increases the instability of cluster memberships and increases the model convergence times, especially during minor drifts. When cluster memberships remain largely stable, global re-clustering is not only excessive but also requires additional steps to warm up the new clusters (e.g., finding optimal cluster model parameters). On the other hand, when many clients experience drifts, moving individual clients can cause substantial shifts in cluster centers and miss the opportunity to create or remove clusters as distribution patterns evolve.
Based on the prior analysis, we opt for a design that starts with client individual migration and triggers global re-clustering selectively (Section 4 explains the detailed design). Our method is visualized as the red curve (“Move Ind.+Global Recluster”) in Figure 1. We retain a low heterogeneity value throughout training and achieve high model accuracy faster (Section  6).
3.2. Re-clustering All Drifted Clients
When we rerun the clustering algorithm every round, we can only easily recluster selected clients (Liu et al., 2023). Since selected clients already get the gradients or model parameters during training, it is easy to compare with the existing clusters and decide whether they experience drifts and need change. In contrast, since unselected clients do not report gradients, we cannot decide whether and where to move them.
However, as mentioned in Section 2.1, pedagogical FL settings typically select only a subset of clients to contribute in each round. When unselected clients experience data drifts and are still stuck in their previous cluster, they might be served by the cluster model trained on data of a different distribution. We plot the difference in clients’ mean test accuracy of re-clustering only selected clients versus all clients that experience data drifts every round in Figure 2(b). Re-clustering all drifted clients offers a 1.5%-4% higher accuracy in most rounds, which is a significant difference given that the final accuracy we achieve on FMoW is 52.4%. This trend demonstrates that only moving selected clients leads to imperfect clusters containing diverging but unselected clients.
Table 1. Heterogeneity of all clients and intra-cluster heterogeneity after gradient-based clustering (lower is better). At the initial stages, gradient-based clustering doesn’t sufficiently decrease heterogeneity due to the model being not well-trained. In contrast, label-based clustering provides a consistent decrease in heterogeneity.
100 1.81 1.82(+0.6%) 1.65(-8.8%)
200 1.80 1.82(+1.1%) 1.64(-8.9%)
500 1.78 1.71(-3.9%) 1.62(-9.0%)
1000 1.74 1.66(-4.6%) 1.56(-10.3%)
1500 1.76 1.63(-7.4%) 1.60(-9.1%)
Global re-clustering based on training outputs incur high overheads. Considering the suboptimal accuracy of only re-clustering the selected clients, our goal is to move all drifted clients promptly. One effective method to learn information from all clients for re-clustering is to train a global model with selected clients and then distribute it to all clients. For example, in FlexCFL, each client performs local training with the global model weights on their own data and reports gradients to the central server, where clustering happens (Duan et al., 2021). The coordinator can then re-cluster all clients according to the gradients.
This approach incurs high communication and computing overheads. For example, in our experiments training ResNet-18 (He et al., 2016) for image classification tasks, devices on average take 116.7 seconds to download the 11 million parameters and upload gradient coordinates. They also need an additional 50.4 seconds to run forward and backward propagation.
3.3. Label-based Instead of Gradient-based Clustering
Another issue with gradient-based re-clustering is that the clustering effectiveness is sensitive to the model quality. Table 1 presents the resulting mean client distance when we perform gradient-based clustering after training a global model for various numbers of rounds. The numbers in parentheses indicate the change in mean client distance relative to that of the global set. A negative value in the parentheses indicates that the generated clusters are overall less heterogeneous than the global set. When we perform gradient-based clustering with the global model at round 100 and round 200, the resulting clusters are more heterogeneous. As we postpone clustering to later rounds, gradient-based clustering achieves an increasingly larger reduction in mean client distance. This trend indicates that gradient-based clustering doesn’t perform well in earlier rounds, when the global model used to collect gradients from all clients hasn’t converged.
We propose using label distribution vectors for clustering (we refer to this approach as “label-based”). Label distribution vectors are available on all clients and enable us to promptly detect any drift by checking for distribution changes. Compared to gradient-based clustering, label-based doesn’t demand parameter or gradient transmission, doesn’t introduce additional computation tasks, and is not sensitive to the timing of client clustering. In Table 1, label-based clustering provides consistent heterogeneity reduction across rounds. Furthermore, as a label distribution vector typically has tens or hundreds of coordinates compared to millions in a full gradient, label-based clustering incurs significantly lower computational overhead. In our experiment of training ResNet-18 on the FMoW dataset, clustering all 302 clients using our label-based solution takes only 0.05 seconds while the gradient-based solution FlexCFL takes 37.92 seconds (note that FlexCFL already accelerates this process through dimensionality reduction using truncated Singular Value Decomposition).
4. Fielding Overview

### v2

# Fielding: Clustered Federated Learning with Data Drift

@@
Abstract
Federated Learning (FL) trains deep models across edge devices without centralizing raw data, preserving user privacy. However, client heterogeneity slows down convergence and limits global model accuracy. Clustered FL (CFL) mitigates this by grouping clients with similar representations and training a separate model for each cluster. In practice, client data evolves over time – a phenomenon we refer to as data drift – which breaks cluster homogeneity and degrades performance. Data drift can take different forms depending on whether changes occur in the output values, the input features, or the relationship between them. We propose Fielding, a CFL framework for handling diverse types of data drift with low overhead. Fielding detects drift at individual clients and performs selective re-clustering to balance cluster quality and model performance, while remaining robust to malicious clients and varying levels of heterogeneity. Experiments show that Fielding improves final model accuracy by 1.9–5.9% and achieves target accuracy 1.16x–2.23x faster than existing state-of-the-art CFL methods.
1 Introduction
@@
Federated Learning (FL) is a distributed machine learning paradigm that enables multiple clients to collaboratively train a shared model while keeping raw data local (Bonawitz et al., 2019; McMahan et al., 2017). FL has gained attention due to its privacy-preserving nature, as it avoids transmitting sensitive data to a central server (Lyu et al., 2024; Huba et al., 2022). This makes it particularly appealing in applications such as mobile text prediction (Yang et al., 2018), energy forecasting (Abdulla et al., 2024), and medical imaging (Zhou et al., 2023).
In practice, FL involves clients with diverse hardware, usage patterns, and data sources, leading to variation in both dataset sizes and distributions: known as data heterogeneity (Li et al., 2022a; Kim et al., 2024). This heterogeneity slows model convergence and degrades model performance (Li et al., 2020a; Zhao et al., 2018; Sattler et al., 2020). Clustered FL (CFL) addresses this by grouping clients with similar data characteristics and training a separate model per cluster (Ghosh et al., 2020; Duan et al., 2021; Liu et al., 2023; Jothimurugesan et al., 2023), improving learning efficiency and accuracy by reducing intra-cluster heterogeneity.
However, real-world deployments face data drifts over time. Data drift typically falls into three categories (Huyen, 2022): (1) label shift, where the output categories frequencies change but the characteristics associated with each category remain stable (Lipton et al., 2018; Garg et al., 2020); (2) covariate shift, where the distribution of the input features changes but the feature-label relationship remains constant (Ganguly and Aggarwal, 2023; Mallick et al., 2022); (3) concept shift, where the underlying input-output relationship changes, rendering a model trained on past data inaccurate on new data (Casado et al., 2022; Jothimurugesan et al., 2023).
As data drift accumulates, intra-cluster heterogeneity increases, and clusters can become as diverse as the full client set (Jothimurugesan et al., 2023; Moreno-Torres et al., 2012). Effective drift handling must address two key challenges: adapting to varying drift magnitudes while maintaining practical efficiency, and supporting diverse drift types.
The first challenge arises when drifts have varying magnitudes. When many clients experience drift, re-clustering based on outdated average weights or gradients becomes invalid, often requiring global re-clustering. FlexCFL (Duan et al., 2021) and IFCA (Ghosh et al., 2020) attempt to adapt incrementally by moving clients one at a time while keeping the number of clusters fixed. They, however, struggle under large-scale shifts as they reassign drifted clients using clusters’ current average gradient or weight, which shift significantly after massive reassignments. Global re-clustering reinitializes all clusters to reflect current data, but incurs significant communication and computation costs: clients might download multiple model replicas and upload new gradients; moreover, the newly formed clusters typically suffer an accuracy dip. Hence, while global re-clustering works for all drift magnitudes, incremental reassignment is more desirable under small drifts for practical reasons. FedDrift (Jothimurugesan et al., 2023) adopts global re-clustering and pays an amplified cost by having each client train on all cluster models. Auxo (Liu et al., 2023) and FedAC (Zhang et al., 2024) reduce overhead by only reassigning and adjusting cluster counts with clients selected for training, but ignore drifted clients that are not selected, reducing overall effectiveness (see Section 2.1).
The second challenge is to detect and adapt to different types of drift. FL frameworks rely on client representations—client-side information used for clustering—to capture drifts. The label distribution vector is a common representation that reflects the label shift. Another is the input embedding: feature vectors derived from a client’s local data using a neural network, then aggregated (e.g., averaging) to represent the input distribution. When the input–output mapping remains stable, input feature shifts often correlate with label distribution shifts so that both representations can produce similar clustering results. However, neither captures changes in the input–output relationship itself – concept drift. Addressing concept drift typically requires loss-based representations, such as loss values, gradients, or gradient directions. For example, Sattler et al. (2021) clusters clients based on gradient directions to identify such shifts. Each representation has trade-offs: Label and embedding vectors capture only marginal distributions, while gradient-based features require additional computation, are sensitive to training noise, and evolve with model updates (Table 3 shows that gradient-based clustering generates better clusters as training progresses and the model becomes more stable). Selecting or combining these representations enables more reliable detection in all types of drift (see Appendix E).
In this work, we present Fielding, a CFL framework that handles multiple types of data drift with low overhead. Fielding makes two key design choices: it combines per-client adjustment with selective global re-clustering, adapting to varying drift magnitudes without incurring the full cost of global re-clustering at every step; and it re-clusters all drifted clients using lightweight, drift-aware representations, keeping client-side computation and communication costs low while enabling efficient information collection from all clients. Our theoretical framework supports all representation choices and provides per-round utility guarantees along with convergence bounds, allowing Fielding to effectively handle all three types of data drift.
To demonstrate practicality, we extended the FedScale engine (Lai et al., 2022) to support streaming data and built a Fielding prototype on top. We evaluated our framework on four image streaming traces with up to 5,078 clients experiencing various drift patterns. In these scenarios, Fielding improves the final accuracy by up to 5.9% and reaches target accuracy as much as 2.23× faster than standard CFL approaches. Moreover, we show that Fielding flexibly accommodates different drift types through pluggable client representations (e.g., label distribution vectors, input embeddings, and gradients), integrates seamlessly with a range of client selection and aggregation schemes, and remains robust against malicious clients and varying degrees of heterogeneity.
2 Fielding
Fielding is a CFL framework that handles multiple types of data drift with low overhead. Managing data drifts effectively requires identifying clients that need reassignment, determining their appropriate clusters, and collecting information to support drift detection and client movement. We begin by motivating the challenges of drift handling, followed by the design decisions that shape Fielding, and then provide an overview of the system.
2.1 Motivation and design choices
We analyze the limitations of existing clustering strategies under real-world drift to motivate the design of Fielding. Our analysis uses the Functional Map of the World (FMoW) dataset (Christie et al., 2018), containing 302 clients (one per unique UTM zone) with time-stamped satellite images labeled by land use (e.g., airport, crop field). Two training rounds correspond to one day. More details on FMoW appear in Section 3. To quantify intra-cluster heterogeneity, we use mean client distance (Lai et al., 2021): for each client, we compute the average pairwise L1 distance between its data distribution and that of clients in the same cluster. We then take the mean across clients.
@@ 2.1 Motivation and design choices
Figure 1: Client heterogeneity of the global set and label distribution-based clusters.
(a) Selective vs. default global re-clustering (positive = selective better).
(b) Global re-clustering vs. Selected clients re-clustering (positive = global better).
Figure 2: Accuracy difference between different re-clustering approaches.
Per-client adjustment fails under many drift. Per-client adjustment methods, such as IFCA (Ghosh et al., 2020) and FlexCFL (Duan et al., 2021), maintain a fixed number of clusters and reassign drifted clients individually based on gradients or model weights. As mentioned in Section 1, such assignments become unstable and lead to poor clustering when many clients drift simultaneously. Figure 1 (“Move Individuals”) highlights this behavior. Between rounds 600–730, a significant drift event increases intra-cluster heterogeneity, and from rounds 740–1340, per-client adjustment fails to restore effective clustering. In fact, heterogeneity exceeds that of the unclustered client set.
Global re-clustering is unstable during small drifts. Global re-clustering resets all client assignments to reflect the current data, and it produces the lowest heterogeneity in our experiments (Figure 1, “Global Recluster”). However, due to the random initialization of k-means (Arthur and Vassilvitskii, 2006; Ahmed et al., 2020), the resulting clusters can vary considerably across rounds, even under small drifts. Combined with cluster warm-up, this leads to test accuracy fluctuations. Figure 2(a) shows the accuracy difference between our selective re-clustering approach (described in Section 2.2) and a baseline that triggers global re-clustering after every drift event. The baseline exhibits unstable accuracy, sometimes matching ours but falling short by up to 5%.
Design choice: combine per-client adjustment with selective global re-clustering. Per-client adjustment breaks in highly dynamic environments where many clients experience drift. Meanwhile, global re-clustering can harm model accuracy by destabilizing cluster assignments and slowing convergence, particularly during minor drifts. To balance these trade-offs, Fielding combines both strategies: starting with individual client migration and triggering global re-clustering only when needed (see Section 2.2 for details). As shown by the red curve (“Move Ind. + Global Recluster”) in Figure 1, this approach maintains low intra-cluster heterogeneity throughout training (see Section 3).
Clustering only selected clients reduces accuracy. Methods such as Auxo (Liu et al., 2023) and FedAC (Zhang et al., 2024) use training outputs—e.g., gradients or model parameters—from selected clients to re-cluster them without any additional cost. However, unselected clients do not provide such information, making it unclear whether they have drifted or where they should be reassigned to. As outlined in Appendix B.1, typical FL settings sample only a subset of clients in each round. When unselected clients drift but remain in their previous clusters, they may receive models trained on data with different distributions, reducing accuracy. Figure 2(b) compares the mean test accuracy when re-clustering only selected clients versus all drifted clients in each round. Re-clustering all drifted clients yields a 1.5%–4% accuracy gain across most rounds—a meaningful improvement given the final FMoW accuracy of 52.4%. These results show that re-clustering only selected clients produces imperfect clusters containing misaligned clients.
Global re-clustering based on training outputs incurs high overheads. Handling all drifted clients requires input from every client, not just those selected for training. Depending on the representation, this method might incur significant overhead. In FlexCFL (Duan et al., 2021), clients train the global model locally and report gradients to the coordinator. In our experiments using ResNet-18 (He et al., 2016) on FMoW, clients spent on average 116.7 seconds for model download and gradient upload, and an additional 50.4 seconds running forward and backward passes on local data.
Design choice: efficient re-clustering of all drifted clients across drift types. To handle drift comprehensively and efficiently, Fielding re-clusters all drifted clients, regardless of whether they were selected for training. Instead of relying on training outputs, it uses lightweight representations that can be collected from all clients with minimal overhead. To support different drift types while maintaining low cost, Fielding allows configurable client representations: it uses label-distribution vectors—small and easy to compute—for detecting label and covariate shifts, and relies on gradients only when needed to capture concept drift. This design avoids accuracy loss caused by stale cluster assignments and eliminates the cost of global model propagation and local retraining.
2.2 System overview
Fielding consists of two core components: a centralized coordinator that manages client representations, performs clustering, and orchestrates model training; and a client-side module that tracks local representations and reports updates when drift is detected. As discussed in Section 2.1, neither per-client adjustment nor global re-clustering alone performs well across all drift scenarios. Fielding adopts a hybrid approach: it uses per-client adjustment by default and triggers global re-clustering only when necessary based on a threshold $\tau$. When clients detect drift, they send updated representations to the coordinator. The coordinator initially reassigns each drifted client to the nearest cluster without updating cluster centers during this phase to ensure deterministic outcomes regardless of client processing order. After all individual adjustments, the coordinator recalculates cluster centers and measures the shift. If any center has moved more than $\tau$, global re-clustering is triggered for all clients; otherwise, the updated cluster membership is finalized. In our prototype, setting $\tau$ to one-third of the average inter-cluster distance worked empirically well. Figure 3 illustrates this clustering process using label-distribution vectors as representation. Clients submit their representations upon registration and whenever drift is detected. The coordinator updates client metadata and assigns each client to the nearest cluster. Before selecting participants for the next round, it checks whether any cluster centers have shifted beyond $\tau$ to determine if global re-clustering is needed.